OpenStudy (anonymous):

Factor this expression completely, then place the factors in the proper location on the grid. 1/8x^3-1/27y^3

OpenStudy (espex):

Assuming your expression is \(\frac{1}{8x^{3}}-\frac{1}{27y^{3}}\) then \(a^3 - b^3 = (a - b) (a^2 + a b + b^2) \)